A complex number z satisfies the equation "z^{2}=3+4i."

Which of the following is true about "z" ?




  1. "(3+4i)^{2}=25+24i"
  2. "z=\pm (2+i)"
  3. "(3+4i)^{2}=9+40i"
  4. "z=\pm (2+2i)"
  5. "(3+4i)^{2}=-7+ 24i"
1
Expert's answer
2021-10-19T16:10:02-0400

Z2=3+4i,Z2-3-4i=0,


using quadratic formula,


Z=(−b±√(b2 − 4ac))/2a


where a=1,b=0,c=(-3-4i)


Z=(-0±√(02-4x1x(-3-4i)))/(2x1)


Z=(0±√(12+16i))/2


Z=(±(4+2i))/2


Z=±(2+i). Hence 2 is correct.




(3+4i)2=(3+4i)(3+4i)


(3+4i)2=9+12i+12i+(4i)2

(3+4i)2=9+24i+16(√(-1))2

(3+4i)2=9+24i+16(-1)


(3+4i)2=9+24i-16


(3+4i)2=(9-16)+24i


(3+4i)2=-7+24i.


Hence 5 is correct.


Therefore 2 and 5 are correct